OpenStudy (anonymous):

graph the line with slope -1/3 passing through the point (1, -4)

OpenStudy (precal):

you could use y=mx+b m=-1/3 that is your slope and x=1 y=-4 solve for b

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{1}{3} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \) get its equation first keep in mind that is a linear equation, thus it makes a line to draw a line all you really need is 2 points so once you have the equation of it, use the "x" and "y" intercepts to graph it